OpenStudy (anonymous):

prove, using the definition of limit that lim(x approaches -2) x^3=-8

OpenStudy (anonymous):

The limit: \[ \lim_{x\to a}f(x) = L \]is defined as: \[ \forall \epsilon>0\;\exists\delta>0\quad\forall x:\;0\lt |x-a|\lt\delta\implies |f(x)-L|\lt \epsilon \]That is to say, For any arbitrary \(\epsilon\) range You must provide a method of finding a \(\delta\) range such that: For any \(x\) is in the \(\delta\) range of \(a\):\[ a-\delta \lt x \lt a+\delta \]Every \(f(x)\) must be within the \(\epsilon\) range of \(L\):\[ L-\epsilon\lt f(x) \lt L+\epsilon \]In general, the method is some function \(g(\epsilon)\) which returns a valid \(\delta\).

OpenStudy (anonymous):

\[ \lim_{x\to (-2)}x^3 = (-8) \]is defined as: \[ \epsilon>0\;\exists\delta>0\quad\forall x:\;0\lt |x- (-2)|\lt\delta\implies |x^3 - (-8)|\lt \epsilon \]

OpenStudy (anonymous):

\[ \begin{array}{rcl} \forall \epsilon>0\;\exists\delta>0\quad\forall x:\; \\ \\ 0\lt |x- (-2)|\lt\delta &\implies& |x^3 - (-8)|\lt \epsilon \\ 0\lt |x + 2|\lt\delta &\implies& |x^3 + 8|\lt \epsilon \\ 0\lt |x + 2|\lt\delta &\implies& |x +2||x^2-2x+ 4|\lt \epsilon \\ \end{array} \]

OpenStudy (anonymous):

One thing to take note is that: \[ |x^2-2x+4| = x^2-2x+4 \]Because it is always positive. Now the trick here is we can't simply say \[ g(\epsilon) = \frac{\epsilon}{x^2-2x+4} \]Our method \(g\) must be independent of \(x\).

OpenStudy (anonymous):

However, one thing we have on our side is that we can let \(\delta\) smaller than it needs to be.

OpenStudy (anonymous):

We can also use the minimum function.

OpenStudy (anonymous):

Suppose we let \(\delta = 1\):\[ \begin{array}{rcccl} 0 &<& |x+2| &<& 1 \\ -1 &<& x+2 &<& 1 \\ -3 &<& x &<& -1 \\ \end{array} \\ \]What happens to \(x^2-2x+4\)?\[ (-3)^2-2(-3)+4 = 19 \\ (-1)^2-2(-1)+4 = 5 \] Remember that we want to minimize \(\delta\) To do so we must to maximize \(x^2-2x+4\), because it is in the denominator. When \(1\) is our delta range, we know that \(19\) is the highest \(x^2-2x+4\) will be. So we let: \[ g(\epsilon) = \frac{\epsilon}{\max(x^2-2x+4)}=\frac{\epsilon}{19} \]There is one issue here though: \(19\) is only the max when \(\delta=1\) and thus \(-3<x<-1\). How about when \(\delta > 1\)? The solution is the minimum function! Never let \(\delta>1\) be the case! No one is forcing us to chose \(\delta > 1\). \[ \delta = g(\epsilon) = \min\left(1,\frac{\epsilon}{19}\right) \] And this concludes the proof.

OpenStudy (anonymous):

@Opcode Tell me what you think.

OpenStudy (anonymous):

Which part doesn't make sense or is hard to follow?

OpenStudy (anonymous):

I probably should have written: \[ |x+2||x^2-2x+4| < \delta |x^2-2x+4| < \epsilon \implies \delta < \frac{\epsilon}{|x^2-2x+4|} \]

OpenStudy (anonymous):

@abb0t Want to double check?

OpenStudy (abb0t):

The last part is how I usually write it. But damn, that's a lot of work! Lol. Nicely done though.

OpenStudy (anonymous):

The tricky part of these proofs is you can't let \(\delta\) be a function of \(x\).

OpenStudy (anonymous):

If you could, then you would be able to prove false things.

OpenStudy (abb0t):

Lol. Let me start from the top.

OpenStudy (abb0t):

Where did the \(x^2+2x-4\) come from?

OpenStudy (anonymous):

\[ x^3+8 = (x+2)(x^2-2x+4) \]

OpenStudy (anonymous):

\[ (a^3+b^3) = (a+b)(a^2-ab+b^2) \]

OpenStudy (abb0t):

I don't think i'm understanding the min/ max portion of it. Because you're only really concerned with what's going on around x = -2 and you already have something familiar. And I think you already have a term that's familiar on the right. If i'm following your steps correctly.

OpenStudy (anonymous):

Well suppose they chose \(\epsilon = 10000\)

OpenStudy (anonymous):

We don't want to choose \(\delta = 10000/19\) because it's too big a range.

OpenStudy (anonymous):

Essentially it is protecting use from big epsilon having us pick too big a delta.

OpenStudy (anonymous):

\[ |x+2| < 10000/19 \implies -10000/19<x < 10000/19-2 \]If we plug \(10000/19\) (which is now allowed) into \(|x^3 + 8| < 10000\) We find out it is false...

OpenStudy (anonymous):

Whoops... \(10000/19-3\) is now allowed, I mean.

OpenStudy (anonymous):

That's why we just want to say \(\delta =1\). =]

OpenStudy (abb0t):

uh...potato?

OpenStudy (anonymous):

The \(\min\) part is really only there for rigor. At some point we need to say "Let's not get any further from \(-2\) than this"

OpenStudy (anonymous):

The \(\min\) is just saying let's got get really far away from \(x\) because to doesn't matter past that point.